OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

Parallel means their slopes are equal then use point slope form y-y1=m(x-x1)

OpenStudy (anonymous):

3x − y = −3 3x + 3 = y slope = 3